Depot report - Saturday 3rd October

It was one of those strange days when not a lot seemed to be visibly happening but when we all collated our notes there was certainly some good progress. Mike did a couple of Inductions with new volunteer Richie Conley and then Tony Kornas from our sponsors Siemens, hopefully Tony will be joining us a volunteer as well.

Mike and Graham had a meeting with Tony to discuss progress on the scheme and discuss ways of presenting it during the current situation with people coming from across the UK. Some good progress was made, and plans formulated. Tony then assisted Graham in installing the new TV for the classroom and this completed the audio-visual system, and all is working.

Bob set about copper greasing and then the cleaning of the PU feet mountings and they are now all in primer and undercoat and will be glossed when the units are fitted. He then moved inside the loco armed with a 'dye' to run down and clean the threads on the floor studs for fitting the feet to. Nuts and washers sourced and added to the studs ready for the units......won't be too long, pictures by Bob Newcombe attached.

Peter V disappeared into 15's No 2 cab to make a start of removing the old cab heater elements and installing the new, he managed to install just under 50% of the new elements, good progress. Little Alex continued his painting on 'ALYCIDON'. In my absence due to the meeting, Geoff made a start on tidying and sorting bits for 15 which had been stored next to the loco but needed to be moved to allow clear access for the return of No 8's cab in about 10 days. After the meeting Mike and Geoff managed to successfully clear all potential obstacles which was great as we thought it may take a couple of days. 

Geoff then moved onto starting to put things together for the refit of No 15's No 1 cab, please see attached pictures by Geoff Brain of his starting point, wait to see the finished article!!. Steve P firstly returned all of 15's engine room pipework he had taken out of store recently to check they were ready for fitting and prepare a shopping list for the replacement rubber hoses.  He also did a similar check with Pete S in the cooler groups, these will all now be ordered from ERIKS. During his tidy up of his workstation he managed to find a box of pipe clamps we thought were missing. He then managed time to drill out some stressed rivets on the 'A' side cantrail ready for Nigel J to refit next week. 

During the course of the afternoon Graham and Pete did some checking on the Cast bogies for 15 to ensure that all was there as work will be starting shortly on the rebuild. They then moved onto checking the Traction Motor cables under 9 which took some of the hit last year. It looks like 8 of the 24 cables will need to be replaced, this includes a couple of the longer ones which go from 'end to end' of the loco......ouch expensive, estimated costs for the cables and fittings is £10,000...if you wish to contribute towards this project please feel free by clicking HERE , alternatively please see where you can download a Standing Order Mandate to join the Friends of the DPS Scheme where you can donate as much as you like either  monthly, quarterly or annually or just make a lump sum donation. A Benefits scheme is being worked on for people who sign up to the Friends scheme.

Words collated by Mike Hallam-Rudd from those present. PU Feet pictures copyright Bob Newcombe and Cab Fitting pictures copyright Geoff Brain.